# Taming the Memory Footprint Crisis: System Design for Production Diffusion LLM Serving

### v1

@@ 1 Introduction
While Autoregressive (AR) LLMs are fundamentally memory-bandwidth bound due to the serial loading of KV states for single-token generation, dLLMs introduce a distinct memory-capacity and compute-intensity bottleneck. By iteratively refining entire sequences via bidirectional attention, dLLMs expose massive intra-sequence parallelism that can saturate GPU compute units. However, this comes at a steep systems cost: the computation of intermediate activations and sequence-wide logits across all tokens creates massive, oscillating activation footprints that shatter the memory assumptions of existing serving engines. Consequently, standard schedulers that optimize for monotonic AR growth fail to manage the dynamic expansion and contraction of dLLM memory demands, leading to severe underutilization or Out-Of-Memory (OOM) crashes. Recent results even demonstrate faster-than-AR inference when dLLM decoding is restructured for block-wise parallelism and cache reuse (e.g., D2F) [wang2025diffusionllmsfasterthanarinference]. Beyond speed, controlled studies indicate that masked/absorbing diffusion can be more data-efficient than size-matched AR models under constrained-data regimes, strengthening the case for dLLM deployment when data, not compute, is the bottleneck [prabhudesai2025diffusionbeatsautoregressivedataconstrained].
@@
Modern AR LLM inference frameworks combine batching, KV caching, and Algorithm-Hardware Co-design to maximize the throughput. However, dLLM inference introduces unique constraints: feature drift and parallel decoding that require specialized handling beyond standard AR designs. First, regarding feature drift, the evolving nature of token representations invalidates append-only caching. To address this, recent works have developed robust validity protocols [ma2025dkvcachecachediffusionlanguage, liu2025dllmcacheacceleratingdiffusionlarge] that selectively refresh or delay cache updates to ensure state consistency without full recomputation. Second, regarding parallel decoding, simultaneously committing multiple tokens risks violating local dependencies. To resolve this, confidence-aware policies [wu2025fastdllmv2efficientblockdiffusion, wang2025diffusionllmsfasterthanarinference] have been proposed to dynamically balance intra-sequence parallelism against generation quality. Together, these contributions have successfully established the algorithmic foundations for correct and efficient dLLM inference.
@@
While the ecosystem for AR LLM serving is robust, featuring mature frameworks [SLED1, vllm, sglang, sheng2023flexgen] like vLLM that optimize throughput and latency, today’s dLLM stacks still lack a comparable end-to-end serving framework. Existing work such as dLLM-Cache [liu2025dllmcacheacceleratingdiffusionlarge], Fast-dLLM [wu2025fastdllmtrainingfreeaccelerationdiffusion, wu2025fastdllmv2efficientblockdiffusion], and Sparse-dLLM [song2025sparsedllmacceleratingdiffusionllms] optimizes isolated components: cache policies, decoding rules, or sparsity patterns. However, no system jointly budgets activation and KV memory under oscillating demand, nor integrates head-level sparsity with a serving scheduler. As a result, existing frameworks either run dLLMs in toy, single-batch settings or encounter OOM and under-utilization when attempting to serve many concurrent requests on production GPUs.
@@ 2.1 Autoregressive LLMs
AR LLM inference proceeds in two phases: prefill, where the model processes the initial prompt and stores the KV cache to prevent redundant recomputation, and decode, where the model generates tokens sequentially by attending to the prefix KV cache. As the sequence grows, the KV cache footprint expands linearly, making memory capacity the primary bottleneck for serving throughput. To address this, modern serving engines employ Continuous Batching and PagedAttention [vllm] to minimize fragmentation and maximize memory utilization. Furthermore, the FlashAttention series [dao2022flashattention, dao2023flashattention2] optimizes the attention kernel itself, enabling fast, memory-efficient exact attention through IO-aware tiling.
@@ 2.2 Diffusion LLMs Inference
However, dLLM inference challenges the established AR LLM inference designs due to the unique constraints of bidirectional attention and iterative denoising. First, the evolving nature of token representations invalidates standard caching assumptions. Unlike autoregressive models where past tokens remain fixed, dLLM tokens are updated at every denoising step. Consequently, “naive” append-only caching which assumes static history leads to feature drift, where stored KV states diverge from the current denoising trajectory, accumulating significant error. To mitigate this, recent systems employ strictly controlled reuse protocols, such as one-step delayed reuse in dKV-Cache [ma2025dkvcachecachediffusionlanguage] or adaptive, feature-aware refresh policies in dLLM-Cache [liu2025dllmcacheacceleratingdiffusionlarge] that selectively recompute volatile states. Second, the ability to commit multiple tokens per step introduces a complex trade-off between parallelism and correctness. While dLLMs can predict parallel blocks, aggressive multi-token commitment can violate local dependencies and degrade coherence. Optimizing this requires confidence-aware policies (e.g., Fast-dLLM [wu2025fastdllmv2efficientblockdiffusion], D2F [wang2025diffusionllmsfasterthanarinference]) that dynamically balance intra-sequence parallelism against generation quality. Finally, the prevalence of mask tokens with low information density renders dLLMs highly amenable to sparsity, motivating dynamic management strategies like attention-guided eviction (Sparse-dLLM [song2025sparsedllmacceleratingdiffusionllms]) or head-specific sparse patterns (SparseD [wang2025sparsedsparseattentiondiffusion]) to concentrate limited GPU capacity on high-importance tokens. Together, these algorithmic adaptations shape unique memory footprints, presenting specific patterns that can be exploited to optimize dLLM serving efficiency.
@@ 2.4 Sparse Attention for dLLMs
Current methods such as Sparse-dLLM [song2025sparsedllmacceleratingdiffusionllms] enforce a single global sparsity pattern to preserve memory contiguity and kernel regularity. Concretely, the system ranks context tokens using an importance score that is aggregated across heads. To jointly capture (i) local neighborhood structure, where relevance may shift within a short window, and (ii) global heavy-hitters that remain consistently useful, these methods apply a local pooling operator with kernel size $w$ to the raw per-head dot-product scores. The resulting sequence-level importance score for context position $j$ is:
@@ 3 Motivation
dKV-Cache [ma2025dkvcachecachediffusionlanguage] Cache Policy Static Request-Level Decoded tokens (delayed) $\approx L\times$ sizeof(Key + Value)
dLLM-Cache [liu2025dllmcacheacceleratingdiffusionlarge] Cache Policy Static Request-Level Prompt + Adaptive Response $\approx L\times$ sizeof(KV + Attn + FFN)
Fast-dLLM [wu2025fastdllmv2efficientblockdiffusion] Inference Framework Static Request-Level Prefix / Suffix Blocks $\approx L\times$ sizeof(Key + Value)
Sparse-dLLM [song2025sparsedllmacceleratingdiffusionllms] Sparsity Algo Static Request-Level Global Top-$k$ Tokens (Shared) $\approx rL\times$ sizeof(Key + Value)
SparseD [wang2025sparsedsparseattentiondiffusion] Sparsity Algo Static Request-Level No Cache; Logical Masks Only None (Stateless)
dInfer [ma2025dinferefficientinferenceframework] Inference Framework Dynamic Request-Level Vicinity Refresh $\approx L\times$ (Key + Value)
@@ 3.1 Static Request-Level Scheduling Wastes Phase Heterogeneity
The Granularity Mismatch. Despite this predictable fluctuation, existing frameworks [ma2025dinferefficientinferenceframework, wu2025fastdllmv2efficientblockdiffusion] typically schedule at the coarse granularity of a request. They treat the internal phase transitions as a black box and must conservatively provision resources for the Refresh phase throughout the request’s entire lifetime. Consequently, during the bandwidth-bound Reuse phases where activation and compute demands drop precipitously, the massive reserved budget sits idle. These systems cannot systematically exploit this “Reuse headroom” to admit additional work, resulting in low average GPU utilization and suboptimal throughput. dLLM-Serve addresses this with a Phase-Multiplexed Greedy Scheduler (§4.4) that schedules at the granularity of individual phases.
@@ 3.2 Unmanaged Peak Activations Constrain Concurrency
Lack of Explicit Budgeting. Current optimizations do not address this “activation boom.” Methods like Fast-dLLM and dKV-Cache [wu2025fastdllmv2efficientblockdiffusion, ma2025dkvcachecachediffusionlanguage] focus on reducing FLOPs but still materialize the full $[B,L,V]$ tensor when they do execute a step.Similarly, frameworks like dInfer [ma2025dinferefficientinferenceframework] optimize execution graphs but leave the logit tensor exposed. Without an explicit mechanism to decompose this tensor, the serving engine must act conservatively which limits the batch size $B$ to accommodate the worst-case allocation of the monolithic projection. This artificial cap on concurrency leaves GPU compute underutilized during the rest of the forward pass. dLLM-Serve explicitly manages this via Logit-Aware Activation Budgeting (§4.3).
@@ 3.3 Logical Sparsity Fails to Reclaim Physical Resources
To further accelerate dLLM inference, recent research has explored sparse attention for dLLMs [song2025sparsedllmacceleratingdiffusionllms]. However, translating theoretical sparsity into system-level gains remains an open challenge due to a disconnect between logical patterns and physical storage.
The Uniformity Trap. Current sparse methods such as Sparse-dLLM [song2025sparsedllmacceleratingdiffusionllms] enforce a uniform mask across all attention heads to simplify memory management. While this preserves memory contiguity, it implicitly optimizes for average relevance. This is detrimental for dLLMs, where different heads often specialize in distinct semantic features. Forcing all heads to attend to the same subset of tokens discards context that is crucial for specific heads, degrading generation quality at aggressive retention ratios.
The Physical Layout Gap. While recent works like SparseD [wang2025sparsedsparseattentiondiffusion] demonstrate that head-specific patterns can restore this quality, they lack a serving-grade implementation. They typically realize sparsity via logical masking. This approach does not save physical memory capacity or reduce bandwidth usage; the system still allocates the full KV buffer and loads irrelevant tokens, only to mask them out during computation. Without a physical data layout that can store irregular, per-head selections contiguously, the theoretical efficiency of sparse attention cannot be realized in a high-throughput serving engine. dLLM-Serve closes this gap with Head-Centric Sparse KV Cache Management (§4.5).
@@ 4.2 Offline Memory Profiler
dLLM-Serve begins with an initialization phase that maps the GPU memory envelope under worst-case serving pressure. Following the methodology used in vLLM [vllm], the profiler executes dummy workloads under the configured max_num_batched_tokens to measure a conservative peak activation footprint for intermediate tensors (e.g., attention/MLP workspaces and GEMM buffers). We add a small guard band to account for allocator fragmentation and kernel selection variability.
@@ 5 Implementation
We implemented dLLM-Serve by extending the open-source Nano-vLLM [nano-vllm] runtime. The implementation adds diffusion-specific control flow including iterative denoising, phase tracking, and sparse KV cache management while reusing the high-performance components already present in Nano-vLLM [nano-vllm], including paged KV cache storage and FlashAttention. The system is implemented in PyTorch and integrates with HuggingFace Transformers.
@@
Nano-vLLM [nano-vllm] is designed for AR LLMs with a single forward pass per decoding step. We extend its ModelRunner into a diffusion-aware runtime that executes an iterative denoising process and exposes the phase information required by our scheduler (§4.4).
@@ 8 Related Work
Inference Optimization for Diffusion LLMs. Recent work translates dLLM parallelism into practical speedups via cross-step reuse, cache-compatible decoding structure, and commitment policies. Cross-step caches selectively reuse KV states between denoising iterations by using delayed or conditioned refresh, similarity-gated updates, or approximate projections, substantially reducing recomputation [ma2025dkvcachecachediffusionlanguage, liu2025dllmcacheacceleratingdiffusionlarge, hu2025flashdlmacceleratingdiffusionlanguage]. Hybrid and blockwise decoders reintroduce cache-friendly structure while preserving within-block bidirectionality through hierarchical caches, parallel pipelines, and adaptive block sizing [wu2025fastdllmv2efficientblockdiffusion, wu2025fastdllmtrainingfreeaccelerationdiffusion, wang2025diffusionllmsfasterthanarinference, lu2025adablockdllmsemanticawarediffusionllm]. Early-commit and verification strategies reduce both the number of steps and the set of active positions, balancing throughput and accuracy [li2025diffusionlanguagemodelsknow, israel2025acceleratingdiffusionllmsadaptive, christopher2025speculativediffusiondecodingaccelerating]. Systems integrations demonstrate end-to-end gains beyond microbenchmarks, and large dLLMs achieve competitive tokens-per-second on modern accelerators [ma2025dinferefficientinferenceframework, song2025seeddiffusionlargescalediffusion].
@@
Cross-step reuse forms the first layer of savings: delayed/conditioned caches (dKV-Cache [ma2025dkvcachecachediffusionlanguage]) and similarity-driven partial refresh (dLLM-Cache [liu2025dllmcacheacceleratingdiffusionlarge]) exploit the empirical stability of KV states between adjacent steps. Approximate projections (FreeCache [hu2025flashdlmacceleratingdiffusionlanguage]) stabilize reuse further when combined with guided unmasking. A second layer targets attention structure itself. SparseD observes head-wise regularities across steps and applies head-specific sparse patterns late in denoising, preserving accuracy while reducing memory traffic at long context [wang2025sparsedsparseattentiondiffusion]. Sparse-dLLM [song2025sparsedllmacceleratingdiffusionllms] concentrates capacity on salient tokens via dynamic cache eviction driven by attention-guided saliency, which lowers both compute and memory footprint as denoising progresses. Complementary mechanisms such as DPad [chen2025dpadefficientdiffusionlanguage] restrict suffix computation with a sliding window and distance-decay dropout, and mask-aware budgeting refines which prompt tokens remain resident per head and layer under tight memory limits [huang2025masktokensprophetfinegrained].

### v2

@@ 1 Introduction
While Autoregressive (AR) LLMs are fundamentally memory-bandwidth bound due to the serial loading of KV states for single-token generation, dLLMs introduce a distinct memory-capacity and compute-intensity bottleneck. By iteratively refining entire sequences via bidirectional attention, dLLMs expose massive intra-sequence parallelism that can saturate GPU compute units. However, this comes at a steep systems cost: the computation of intermediate activations and sequence-wide logits across all tokens creates massive, oscillating activation footprints that shatter the memory assumptions of existing serving engines. Consequently, standard schedulers that optimize for monotonic AR growth fail to manage the dynamic expansion and contraction of dLLM memory demands, leading to severe underutilization or Out-Of-Memory (OOM) crashes. Recent results even demonstrate faster-than-AR inference when dLLM decoding is restructured for block-wise parallelism and cache reuse (e.g., D2F) [20]. Beyond speed, controlled studies indicate that masked/absorbing diffusion can be more data-efficient than size-matched AR models under constrained-data regimes, strengthening the case for dLLM deployment when data, not compute, is the bottleneck [16].
@@
Modern AR LLM inference frameworks combine batching, KV caching, and Algorithm-Hardware Co-design to maximize the throughput. However, dLLM inference introduces unique constraints: feature drift and parallel decoding that require specialized handling beyond standard AR designs. First, regarding feature drift, the evolving nature of token representations invalidates append-only caching. To address this, recent works have developed robust validity protocols [14, 12] that selectively refresh or delay cache updates to ensure state consistency without full recomputation. Second, regarding parallel decoding, simultaneously committing multiple tokens risks violating local dependencies. To resolve this, confidence-aware policies [22, 20] have been proposed to dynamically balance intra-sequence parallelism against generation quality. Together, these contributions have successfully established the algorithmic foundations for correct and efficient dLLM inference.
@@
While the ecosystem for AR LLM serving is robust, featuring mature frameworks [11, 9, 24, 17] like vLLM that optimize throughput and latency, today’s dLLM stacks still lack a comparable end-to-end serving framework. Existing work such as dLLM-Cache [12], Fast-dLLM [23, 22], and Sparse-dLLM [18] optimizes isolated components: cache policies, decoding rules, or sparsity patterns. However, no system jointly budgets activation and KV memory under oscillating demand, nor integrates head-level sparsity with a serving scheduler. As a result, existing frameworks either run dLLMs in toy, single-batch settings or encounter OOM and under-utilization when attempting to serve many concurrent requests on production GPUs.
@@ 2.1 Autoregressive LLMs
AR LLM inference proceeds in two phases: prefill, where the model processes the initial prompt and stores the KV cache to prevent redundant recomputation, and decode, where the model generates tokens sequentially by attending to the prefix KV cache. As the sequence grows, the KV cache footprint expands linearly, making memory capacity the primary bottleneck for serving throughput. To address this, modern serving engines employ Continuous Batching and PagedAttention [9] to minimize fragmentation and maximize memory utilization. Furthermore, the FlashAttention series [3, 4] optimizes the attention kernel itself, enabling fast, memory-efficient exact attention through IO-aware tiling.
@@ 2.2 Diffusion LLMs Inference
However, dLLM inference challenges the established AR LLM inference designs due to the unique constraints of bidirectional attention and iterative denoising. First, the evolving nature of token representations invalidates standard caching assumptions. Unlike autoregressive models where past tokens remain fixed, dLLM tokens are updated at every denoising step. Consequently, “naive” append-only caching which assumes static history leads to feature drift, where stored KV states diverge from the current denoising trajectory, accumulating significant error. To mitigate this, recent systems employ strictly controlled reuse protocols, such as one-step delayed reuse in dKV-Cache [14] or adaptive, feature-aware refresh policies in dLLM-Cache [12] that selectively recompute volatile states. Second, the ability to commit multiple tokens per step introduces a complex trade-off between parallelism and correctness. While dLLMs can predict parallel blocks, aggressive multi-token commitment can violate local dependencies and degrade coherence. Optimizing this requires confidence-aware policies (e.g., Fast-dLLM [22], D2F [20]) that dynamically balance intra-sequence parallelism against generation quality. Finally, the prevalence of mask tokens with low information density renders dLLMs highly amenable to sparsity, motivating dynamic management strategies like attention-guided eviction (Sparse-dLLM [18]) or head-specific sparse patterns (SparseD [21]) to concentrate limited GPU capacity on high-importance tokens. Together, these algorithmic adaptations shape unique memory footprints, presenting specific patterns that can be exploited to optimize dLLM serving efficiency.
@@ 2.4 Sparse Attention for dLLMs
Current methods such as Sparse-dLLM [18] enforce a single global sparsity pattern to preserve memory contiguity and kernel regularity. Concretely, the system ranks context tokens using an importance score that is aggregated across heads. To jointly capture (i) local neighborhood structure, where relevance may shift within a short window, and (ii) global heavy-hitters that remain consistently useful, these methods apply a local pooling operator with kernel size $w$ to the raw per-head dot-product scores. The resulting sequence-level importance score for context position $j$ is:
@@ 3 Motivation
dKV-Cache [14] Cache Policy Static Request-Level Decoded tokens (delayed) $\approx L\times$ sizeof(Key + Value)
dLLM-Cache [12] Cache Policy Static Request-Level Prompt + Adaptive Response $\approx L\times$ sizeof(KV + Attn + FFN)
Fast-dLLM [22] Inference Framework Static Request-Level Prefix / Suffix Blocks $\approx L\times$ sizeof(Key + Value)
Sparse-dLLM [18] Sparsity Algo Static Request-Level Global Top-$k$ Tokens (Shared) $\approx rL\times$ sizeof(Key + Value)
SparseD [21] Sparsity Algo Static Request-Level No Cache; Logical Masks Only None (Stateless)
dInfer [15] Inference Framework Dynamic Request-Level Vicinity Refresh $\approx L\times$ (Key + Value)
@@ 3.1 Static Request-Level Scheduling Wastes Phase Heterogeneity
The Granularity Mismatch. Despite this predictable fluctuation, existing frameworks [15, 22] typically schedule at the coarse granularity of a request. They treat the internal phase transitions as a black box and must conservatively provision resources for the Refresh phase throughout the request’s entire lifetime. Consequently, during the bandwidth-bound Reuse phases where activation and compute demands drop precipitously, the massive reserved budget sits idle. These systems cannot systematically exploit this “Reuse headroom” to admit additional work, resulting in low average GPU utilization and suboptimal throughput. dLLM-Serve addresses this with a Phase-Multiplexed Greedy Scheduler (§4.4) that schedules at the granularity of individual phases.
@@ 3.2 Unmanaged Peak Activations Constrain Concurrency
Lack of Explicit Budgeting. Current optimizations do not address this “activation boom.” Methods like Fast-dLLM and dKV-Cache [22, 14] focus on reducing FLOPs but still materialize the full $[B,L,V]$ tensor when they do execute a step.Similarly, frameworks like dInfer [15] optimize execution graphs but leave the logit tensor exposed. Without an explicit mechanism to decompose this tensor, the serving engine must act conservatively which limits the batch size $B$ to accommodate the worst-case allocation of the monolithic projection. This artificial cap on concurrency leaves GPU compute underutilized during the rest of the forward pass. dLLM-Serve explicitly manages this via Logit-Aware Activation Budgeting (§4.3).
@@ 3.3 Logical Sparsity Fails to Reclaim Physical Resources
To further accelerate dLLM inference, recent research has explored sparse attention for dLLMs [18]. However, translating theoretical sparsity into system-level gains remains an open challenge due to a disconnect between logical patterns and physical storage.
The Uniformity Trap. Current sparse methods such as Sparse-dLLM [18] enforce a uniform mask across all attention heads to simplify memory management. While this preserves memory contiguity, it implicitly optimizes for average relevance. This is detrimental for dLLMs, where different heads often specialize in distinct semantic features. Forcing all heads to attend to the same subset of tokens discards context that is crucial for specific heads, degrading generation quality at aggressive retention ratios.
The Physical Layout Gap. While recent works like SparseD [21] demonstrate that head-specific patterns can restore this quality, they lack a serving-grade implementation. They typically realize sparsity via logical masking. This approach does not save physical memory capacity or reduce bandwidth usage; the system still allocates the full KV buffer and loads irrelevant tokens, only to mask them out during computation. Without a physical data layout that can store irregular, per-head selections contiguously, the theoretical efficiency of sparse attention cannot be realized in a high-throughput serving engine. dLLM-Serve closes this gap with Head-Centric Sparse KV Cache Management (§4.5).
@@ 4.2 Offline Memory Profiler
dLLM-Serve begins with an initialization phase that maps the GPU memory envelope under worst-case serving pressure. Following the methodology used in vLLM [9], the profiler executes dummy workloads under the configured max_num_batched_tokens to measure a conservative peak activation footprint for intermediate tensors (e.g., attention/MLP workspaces and GEMM buffers). We add a small guard band to account for allocator fragmentation and kernel selection variability.
@@ 5 Implementation
We implemented dLLM-Serve by extending the open-source Nano-vLLM [5] runtime. The implementation adds diffusion-specific control flow including iterative denoising, phase tracking, and sparse KV cache management while reusing the high-performance components already present in Nano-vLLM [5], including paged KV cache storage and FlashAttention. The system is implemented in PyTorch and integrates with HuggingFace Transformers.
@@
Nano-vLLM [5] is designed for AR LLMs with a single forward pass per decoding step. We extend its ModelRunner into a diffusion-aware runtime that executes an iterative denoising process and exposes the phase information required by our scheduler (§4.4).
@@ 8 Related Work
Inference Optimization for Diffusion LLMs. Recent work translates dLLM parallelism into practical speedups via cross-step reuse, cache-compatible decoding structure, and commitment policies. Cross-step caches selectively reuse KV states between denoising iterations by using delayed or conditioned refresh, similarity-gated updates, or approximate projections, substantially reducing recomputation [14, 12, 6]. Hybrid and blockwise decoders reintroduce cache-friendly structure while preserving within-block bidirectionality through hierarchical caches, parallel pipelines, and adaptive block sizing [22, 23, 20, 13]. Early-commit and verification strategies reduce both the number of steps and the set of active positions, balancing throughput and accuracy [10, 8, 2]. Systems integrations demonstrate end-to-end gains beyond microbenchmarks, and large dLLMs achieve competitive tokens-per-second on modern accelerators [15, 19].
@@
Cross-step reuse forms the first layer of savings: delayed/conditioned caches (dKV-Cache [14]) and similarity-driven partial refresh (dLLM-Cache [12]) exploit the empirical stability of KV states between adjacent steps. Approximate projections (FreeCache [6]) stabilize reuse further when combined with guided unmasking. A second layer targets attention structure itself. SparseD observes head-wise regularities across steps and applies head-specific sparse patterns late in denoising, preserving accuracy while reducing memory traffic at long context [21]. Sparse-dLLM [18] concentrates capacity on salient tokens via dynamic cache eviction driven by attention-guided saliency, which lowers both compute and memory footprint as denoising progresses. Complementary mechanisms such as DPad [1] restrict suffix computation with a sliding window and distance-decay dropout, and mask-aware budgeting refines which prompt tokens remain resident per head and layer under tight memory limits [7].